OpenStudy (anonymous):

PLEASE HEEEELP Write the equation in slope-intercept form. Then find the slope and y-intercept of each line. -1/2x - y = 3/4

OpenStudy (hlambach):

Slope intercept form is y=mx+b, m being the slope and b the y-intercept.

OpenStudy (anonymous):

Thanks, could you help me plug in the equation? @HelpBlahBlahBlah @Kayleedid @linn99123 @SithsAndGiggles @rose21 @paki

OpenStudy (anonymous):

@hlambach?

OpenStudy (hlambach):

You don't need to plug it it, you just need to isolate y.

OpenStudy (anonymous):

Okay how do I do that

OpenStudy (hlambach):

So get -1/2x to the other side and you have the answer.

OpenStudy (anonymous):

for finding the slope and y-intercept? what about changing the equation to slope intercept form

OpenStudy (hlambach):

if you get -1/2x to the other side, it IS in slope-intercept form.

OpenStudy (hlambach):

Trust me, just do it. :)

OpenStudy (anonymous):

then what will be the slope and y-intercept

OpenStudy (hlambach):

isolate y, meaning it should look like y=........

OpenStudy (hlambach):

Slope intercept form is y=mx+b, m being the slope and b the y-intercept.

OpenStudy (hlambach):

So the number before x will be the slope, and the number added will be the y-intercept. But first, isolate y.

OpenStudy (anonymous):

please could you show me the steps? like after putting -1/2x what do i do with it could you show me

OpenStudy (hlambach):

OKay, to get anything to the other side, you have to do the opposite. Since it is -1/2x, meaning it is subtracting 1/2x, what is the opposite of subtracting?

OpenStudy (anonymous):

adding

OpenStudy (anonymous):

so i have to add -1/2x?

OpenStudy (hlambach):

Yep, so you ADD 1/2x to both sides. -1/2x-y=3/4 +1/2x +1/2x __________________ -1/2x +1/2x equals what?

OpenStudy (anonymous):

uhhhh 0?

OpenStudy (hlambach):

Yep. But you also add it to the other side of the equal sign, so it would be -y=1/2x+3/4

OpenStudy (hlambach):

Get how I got there?

OpenStudy (anonymous):

that is the slope form?

OpenStudy (hlambach):

No, because that form is y=mx+b the one we have is -y=1/2x+3/4 You need to get rid of the negative.

OpenStudy (anonymous):

oh.. by adding?

OpenStudy (hlambach):

This time no. You don't want to get rid of the y, just the negative sign. You want to make y positive.

OpenStudy (anonymous):

oh okay so how do we do that

OpenStudy (hlambach):

Negative times what equals positive?

OpenStudy (anonymous):

umm positive? i dont know

OpenStudy (hlambach):

Yep.

OpenStudy (anonymous):

okay can you show me

OpenStudy (hlambach):

When there is no number in front of a variable, it is understood that it is 1. Ex: x = 1x In this case, since it is negative, it is -1. so -y=-1y Since you are multiplying y by -1, you need to do the opposite, which is divide by -1. This "/" means divide. -y = 1/2x+3/4 /-1 /-1 Now you divide everything on the other side by -1.

OpenStudy (anonymous):

okay what would be the final result?

OpenStudy (hlambach):

Well, what do you think?

OpenStudy (anonymous):

um.. i dont know :/

OpenStudy (hlambach):

When you divide something by -1, everything becomes a negative. So: y=-1/2x-3/4 Get it?

OpenStudy (anonymous):

yeah, so that is the slope form?

OpenStudy (hlambach):

Yes. Now what is the slope and y intercept?

OpenStudy (anonymous):

y intercept would be -1/2x? im not sure

OpenStudy (hlambach):

Remember that in slope intercept form, (y=mx+b) m, which is the number before the x, is the slope, and b is the y-intercept. so in y=-1/2x-3/4

OpenStudy (hlambach):

Which numbers are m and b?

OpenStudy (anonymous):

-1/2 and -3/4?

OpenStudy (hlambach):

Yep!

OpenStudy (anonymous):

Thank you! could you please help with another one? please!

OpenStudy (anonymous):

I really need help!

OpenStudy (hlambach):

SUre, if I can.

OpenStudy (anonymous):

Thanks so much, I will tag you in another post so you can get more medals!

OpenStudy (hlambach):

Okay, thanks!